A salesman is known to sell a product in 3 out of 5 attempts. while another salesman in 2 out of 5 attempts. Find the probability that
1.no sale while happen
2. either of them will succeed in selling the product
1
Expert's answer
2013-04-26T09:33:28-0400
1. First salesman is known to sell a product in 3 out of 5attempts, so the probability he'll make a sale is 3/5 and hence the probability
he'll not make a sale is (1-3/5). In the same way the probability that second
salesman will not make a sale is (1-2/5). Since this two events are
independent, the probability that no sale will happen is (1-3/5)*(1-2/5)=0.24
according to product rule.
2. The probability either of them will succeed = 1-Probability (No sale will
happen) = 1 - 0.24 = 0.76